# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v5

@@ 4 Conclusion and Limitation
We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation, along with a self-collected dataset for novel view synthesis under multi-view long-view scenarios. Our proposed method splits the input multi-view videos into different chunks, trains different model branches, and performs reconstruction accordingly. Each branch encodes spatial and temporal features separately using the multi-resolution hash encoding structure. It leverages the redundancy of the dynamic scenes and reuses the spatial features in the first branch to achieve fast convergence, high fidelity, and superior scalability. On the prevailing DyNeRF dataset, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit offline baselines, achieves comparable quality, and maintains the speed advantage over pure-MLP-based offline methods. Compared with online baselines, it delivers state-of-the-art quality and achieves a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost. On the long video dataset, our method delivers high quality and achieves fast convergence with only a fraction of the storage cost. However, like most NeRF-based methods, it still requires querying the field function and is not suitable for real-time rendering. Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning, enabling real-time rendering in long video sequences for wide applications.
@@ Appendix A Per-scene results
Table 12: Detailed results of the proposed CD-NGP and other hash-based methods on our long video dataset.

### v6

@@ 4 Conclusion and Limitation
We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation, along with a self-collected dataset for novel view synthesis under multi-view long-view scenarios. Our approach partitions the input multi-view videos into different chunks and uses individual model branches to reconstruct the scene accordingly. Each branch encodes spatial and temporal features separately using the multi-resolution hash encoding structure. CD-NGP leverages the redundancy of the dynamic scenes and reuses the spatial features in the first branch to achieve fast convergence, high fidelity, and superior scalability. On the prevailing DyNeRF dataset, it consumes much less memory, achieves comparable quality with a smaller model size than most recent explicit offline baselines, and maintains the speed advantage over traditional pure MLP-based offline methods. Compared with online baselines, it delivers state-of-the-art quality and achieves a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost. On the long video dataset, our method delivers high quality and achieves fast convergence with only a fraction of the storage cost. However, like most NeRF-based methods, it still requires querying the field function and is not suitable for real-time rendering. Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning, enabling real-time rendering in long video sequences for wide applications.
@@ Appendix A Per-scene results
Table 13: Detailed results of the proposed CD-NGP and other methods on our long video dataset.
@@
Appendix B Access to the long video dataset.
We provide a public link to access our self-collected long video dataset. It can be accessed via the following BaiduNetdisk link:
Appendix C Discussion on rendering latency.
The proposed method CD-NGP fuses spatial and temporal features in a multi-branch NeRF model, which is efficient in terms of training and memory consumption. Moreover, the feature combination in the latent space avoids querying multiple fields for rendering, and thus achieves better rendering latency. However, the volume rendering process adopted from NeRF-variants is inherently computationally intensive, which limits the real-time rendering capability. In this section, we provide a detailed analysis of the rendering latency of our method on the DyNeRF dataset. Taking the cut roasted beef scene in the DyNeRF dataset (at $1352\times 1014$ resolution) as an example, we provide a detailed analysis of the rendering latency in Table 14. The rendering process of CD-NGP is divided into three parts: ray marching, NeRF model, and pixel-level volume rendering as follows:
  • Ray marching: The ray marching process is the same as that in the implementation [25] of Instant-NGP [34], which is a standard process in hash-based NeRF. This process generates a split (batch) of sampling coordinates from the camera to the scene along each ray, and each image requires multiple passes to render. Typically, each pass generates a split of up to 9M samples on the 1M rays (1M pixels), and rendering the entire image requires around 200M samples.
  • NeRF model: As elaborated in the manuscript, our CD-NGP is a multi-branch NeRF model that incorporates hash-based encoders and MLP-based decoders. As shown in Table 14, both the spatial and temporal hash encoders are efficient and lightweight, requiring only 7 ms and 5 ms, respectively. The primary computational cost arises from the two MLP decoders (55 ms in total), which is common and inevitable in NeRF-based methods. Although CD-NGP avoids multiple passes of MLP decoders by fusing features in the latent space, one fused decoding pass of MLP decoders is still required for each sampled point. This fused pass takes approximately 33 ms for the XYZT-fusion MLP and 22 ms for the RGB MLP, resulting in a total NeRF model latency of around 76 ms per pass.
  • Volume rendering: This process loads the optical attributes $\sigma,\textbf{c}$ of the sampled points and computes the pixel color C by numerical integral (accumulating the optical attributes along the ray). The rendering process is similar to that of Instant-NGP [34] and other NeRF-based methods [25]. If the accumulated opacity of a ray (pixel) reaches $10^{-4}$, the sampling and rendering are terminated. Thus, the whole rendering time is smaller than the product of the number of passes and the time of each pass.
Table 14: Rendering time breakdown and point scale of each component.
Module Time Scale
Ray marching 3ms $\sim$1M pixels
NeRF model
    Spatial representation 7ms $\sim$9M samples
    Time representation 5ms $\sim$9M samples
    XYZT-fusion MLP 33ms $\sim$9M samples
    RGB MLP 22ms $\sim$9M samples
    Total (NeRF model) 76ms $\sim$9M samples
Rendering a split 2ms $\sim$1M pixels
Rendering the whole image 500ms $\sim$200M samples
As illustrated, each query of the hash table-based NeRF model takes less than 100ms, but the pixel-based volume rendering process requires multiple passes to render the whole image. This limitation is a common issue in all NeRF-based methods and is orthogonal to the contribution of the proposed feature fusion method and continual learning pipeline.
Appendix D Comparison with real-time processing method.
We compare the per-scene results of our method and MVSGS-FT on the DyNeRF dataset in Table 15. Our method produces comparable quality with MVSGS-FT in the first 4 scenes and is more robust in the last sear steak scene. The method MVSGS-D does not support the standard $1352\times 1014$ resolution on the DyNeRF dataset and produces limited quality as shown in Fig. 11, we thus exclude it from Table 15.
Refer to caption Refer to caption
(a) (b)
Figure 10: Output images of MVSGS-FT on the DyNeRF dataset. (a) iteration 3,000, (b) iteration 5,000.
Figure 11: Output images of MVSGS-D on the DyNeRF dataset. The left side is the ground truth, and the right side is the rendering output.
Table 15: Detailed results of the proposed CD-NGP and real-time processing method MVSGS.
Model Flame Salmon Cook Spinach Cut Roasted Beef
[…]
CD-NGP 26.37 0.046 0.289 30.76 0.024 0.179 31.03 0.024 0.180
MVSGS-FT 24.75 0.054 0.190 30.57 0.023 0.123 29.68 0.025 0.131
Model Flame Steak Sear Steak Average
[…]
CD-NGP 31.27 0.021 0.174 31.71 0.019 0.170 30.23 0.027 0.198
MVSGS-FT 30.02 0.027 0.126 15.26 0.301 0.722 26.05 0.086 0.258
We provide the rendered images of MVSGS-FT for better visualization in Fig. 10. We note that there are two approaches for fine-tuning the MVSGS: the first is to use the output of the MVS network to re-initialize Gaussians for each frame, and the second is to use the Gaussians of the previous frame for initialization. However, we find that the second approach leads to unstable training and produces unsatisfactory results (much lower PSNRs). Therefore, we only report the metrics of the first approach and refer to it as MVSGS-FT.
Appendix E Implementation Details about the spatial and temporal hash encoding
E.1 Implementation
In the implementation of tiny-cuda-nn, the authors have provided interfaces for hash-encoders with different input dimensions, such as 1D, 2D, and 3D. However, the 1D hash encoding is not enabled by default. In our implementation, we enable the interfaces for one-dimension hash encodings of tiny-cuda-nn [33] to support the temporal encoding in our continual learning framework. Specifically, we use the 1D hash encoding for the time dimension and the 3D hash encoding for the spatial dimensions.
E.2 Scalability on size.
We measure the hash encoder size and the inference latency of different configurations of the encodings in Table 16. The latency is measured on a single NVIDIA RTX 3090 GPU. Since all hash tables are stored in float32 precision, each element of the table consumes $F$ parameters, and the hash encoder contains $L$ hash tables for different encoding resolutions. The upper bound of the total hash encoder size is $4\times L\times F\times 2^{P}$ bytes. Since some spatial hash table resolutions are coarse (e.g., $16\times 16\times 16=2^{12}<2^{14}<2^{P_{1}=19}$), and the tables are not fully filled, the actual hash encoder size is smaller than the upper bound. Clearly, the hash encoder size is primarily influenced by the length of the hash tables (i.e., $2^{P}$). By controlling the length of the hash tables and reusing features from the base spatial hash table, we can achieve a good trade-off between hash encoder size and rendering quality.
E.3 Scalability on inference speed.
Table 16 additionally presents the inference speed across different configurations. We render the first frame of the cut roasted beef scene from the DyNeRF dataset, which has a resolution of $1352\times 1014$. The latency increases approximately linearly with both the number of layers $L$ and the feature vector length $F$, with $L$ exhibiting a greater impact. Increasing the hash table length $P$ also affects inference speed, primarily due to reduced memory locality for larger table sizes.
Table 16: Inference speed and model size under different configurations. L: levels, F: feature dim, P: log2 values of the hash table length. Base and auxiliary denote the hash encoders of the base spatial representation and auxiliary spatial representation, respectively. Temporal denotes the hash encoder of the time dimension.
Type L F P Size (MB) Samples Latency(ms)
base 12 2 19 37.83 5483712 3.33
auxiliary 12 2 14 1.41 5483712 3.16
temporal 2 20 7 0.006 5483712 2.89
base 16 2 19 51.03 5483712 4.48
auxiliary 16 2 14 1.91 5483712 4.27
temporal 2 20 7 0.006 5483712 2.96
base 24 2 19 77.53 5483712 6.82
auxiliary 24 2 14 2.86 5483712 6.51
temporal 2 20 7 0.006 5483712 2.89
base 12 4 19 75.65 5483712 5.33
auxiliary 12 4 14 2.82 5483712 5.08
temporal 2 20 7 0.006 5483712 3.01
base 12 2 21 140.83 5483712 3.64
auxiliary 12 2 14 1.41 5483712 3.16
temporal 2 20 7 0.006 5483712 3.19
base 12 2 20 73.82 5483712 3.45
auxiliary 12 2 14 1.41 5483712 3.16
temporal 2 20 7 0.006 5483712 3.23
base 12 2 19 37.83 5483712 3.31
auxiliary 12 2 16 5.26 5483712 3.24
temporal 2 20 7 0.006 5483712 2.87
base 12 2 19 37.83 5483712 3.30
auxiliary 12 2 14 1.41 5483712 3.14
temporal 1 20 7 0.003 5483712 1.39
base 12 2 19 37.83 5483712 3.29
auxiliary 12 2 14 1.41 5483712 3.15
temporal 1 40 7 0.006 5483712 2.87